A survey of 144 retail stores revealed that the average price of a microwave was $375 with a standard error of $20

What is the 99% confidence interval to estimate the true cost of the microwave?

Answer by stanbon(75887)

A survey of 144 retail stores revealed that the average price of a microwave was $375 with a standard error of $20
What is the 99% confidence interval to estimate the true cost of the microwave?
------------------
x-bar = 375
---
ME = invT(0.995 with df=143)*SE
ME = (2.6106)*20 = 52.21

-----------
99% CI: 375-52.21 < u < 375+52.21
====================================
Cheers,
Stan H.
==================